OpenStudy (anonymous):

the sum of twice a number and 14 less than the number is the same as the difference -34 and the number. what is the number?

OpenStudy (anonymous):

let the number be x 2x + (x-14) = hmm what exactly dp u mean by difference -34 and the number?

OpenStudy (anonymous):

do u mean the number - 34?

OpenStudy (anonymous):

Let the number be x then ATQ, 2x + (x-14) = -34 - x this gives 3x -14 = -34 -x 3x + x = -34 + 14 4x = - 20 x = -5 So the number is -5

OpenStudy (anonymous):

I hv understood it as "the difference of -34 and the number" and solved it. I think Bridget missed the word "of" before -34

OpenStudy (anonymous):

u right thanku

OpenStudy (anonymous):

"Good Answer" button click pls